# Probing Binary Lens Caustics with Gravitational Waves: A Uniform Approximation Approach

### v2

@@ V.2 Amplification
$|F_{TA}|^{*}=\frac{1}{\sqrt{\rho_{c}\,|T_{11}|}}{\cal A}(u^{*})=\sqrt{2/|T_{11}|}(\rho_{c}\,d_{c})^{-1/4},$ (32)
@@
In the last line of this equation, the next-order term is included [68]. Since the dominant term in both approximations agrees, it can be reliably used to estimate the peak amplification. Given that GW detectors measure strain—sensitive to both amplitude and phase rather than intensity—the appropriate amplification quantity is the linear factor $\mu_{\mathrm{GW}}=|F|$. This leads to the expression:
$\mu_{\mathrm{GW}}=\sqrt{2/|T_{11}|}(\rho_{c}\,d_{c})^{-1/4}\equiv\mu_{0}\,d_{c}^{-1/4}$ (34)
where the prefactor $\mu_{0}(b)$ depends only on the binary separation and is therefore fixed for a given lens mass distribution. In contrast, the final term, determined by the caustic width $d_{c}$, captures the frequency (or wavelength) dependence of the incoming wave. For EM waves, where the observations are based on intensity, the relevant quantity is the magnification $\mu_{\mathrm{EM}}=|F|^{2}=\mu_{0}^{2}\,d_{c}^{-1/2}$.
Therefore, the peak amplification is governed by the characteristic scales of the caustic. Using Eq. (31), we find that it scales with the wavelength as $\sim\lambda^{-1/6}$, and with the lens mass as $\sim M_{Lz}^{1/6}$ as shown in Fig. 4. This leads to the following expression:
$\mu_{\mathrm{GW}}\simeq 2.5\ |T_{11}|^{-1/2}\,\rho_{c}^{-1/3}\left(\frac{\lambda}{10^{6}{\rm m}}\right)^{-1/6}\left(\frac{M_{Lz}}{10^{3}M_{\odot}}\right)^{1/6}$ (35)
@@
As an example, consider a binary lens system with $b=0.7$ and a frequency of 500 Hz. In this case, we find that the maximum amplification is $\mu_{\mathrm{GW}}\simeq 2$ for a lens mass of $M=5\times 10^{3}M_{\odot}$, and $\mu_{\mathrm{GW}}\simeq 22.3$ for $M=10^{10}M_{\odot}$. For comparison, the corresponding magnification values for electromagnetic (EM) waves will be the square of these values.
@@ VI Conclusions
This study introduces a novel approach to understanding GW diffraction near caustics created by a binary lens system. The Uniform Approximation proves to be a robust method for modeling the GW field near the fold, addressing the limitations of both Geometrical Optics and Transitional Asymptotics methods. By ensuring continuity across the caustic, the UA effectively links wave behavior near the fold to the far-field GO description, providing a comprehensive solution for wave propagation in caustic regions.
@@ Appendix A Images and caustics of a binary lens
where $\mathbf{x_{A}}$ and $\mathbf{x_{B}}$ are the positions of the two lensing masses on the lens plane, and $\gamma_{1}$ and $\gamma_{2}$ are the reduced masses (see Section II). If the $x_{1}$ axis aligns with the line connecting the masses and the origin is at its midpoint, the mass coordinates satisfy $\mathbf{x_{A}}=-\mathbf{x_{B}}=\mathbf{x_{L}}=(b,0)$.
Figure 5: Source plane with caustics (left panel) and lens plane with critical lines (right panel) for a binary lens system with equal masses and $b=0.7$. A source is placed at two different positions, both at a distance of $|\tilde{y}_{2}|=0.1$ from the fold: one inside the caustic (red star) and one outside (black star). Their corresponding images in the lens plane are marked with dots of the same color. The location of the fold point is indicated by a black circle, while the point of coalescence of the two stationary points in the lens plane is denoted by a black triangle.
@@
Figure 8: Absolute value of the transmission factor computed with UA as a function of the frequency and the separation of the source with respect to the fold. The separation between the lenses is fixed to $b=0.7$ and the position of the fold is shown by a black line. The white continuous and dashed lines indicate the maxima computed with UA (40) and TA (41), respectively.
@@ Appendix B Interference Fringes
where $D=5\rho_{c}^{-1}\left[3/2+(2\rho_{c}^{\prime}/3)^{2}\right]^{-1}$. By substituting the phase difference by Eq. (39), we obtain the source positions corresponding to the extrema of the transmission factor. The dependence of these positions on the caustic curvature and the dimensionless frequency is nontrivial, resulting in a complex scaling for the fringe separation. Fig. 8 shows a colormap of the absolute value of the transmission factor as a function of the source’s distance from the fold and the dimensionless frequency, where Eq. (40) is used to visualize the maxima and minima.

### v3

@@ V.2 Amplification
$|F_{\rm TA}|^{*}=\frac{1}{\sqrt{\rho_{c}\,|T_{11}|}}{\cal A}(u^{*})=\sqrt{2/|T_{11}|}(\rho_{c}\,d_{c})^{-1/4},$ (32)
@@
In the last line of this equation, the next-order term is included [72]. Since the dominant term in both approximations agrees, it can be reliably used to estimate the peak amplification. Given that GW detectors measure strain—sensitive to both amplitude and phase rather than intensity—the strain signal near the fold is amplified by a factor:
$|F|_{\rm max}=F_{0}\,d_{c}^{-1/4}\quad\text{with}\quad F_{0}=\sqrt{2/|T_{11}|}\,\rho_{c}^{-1/4}$ (34)
where the prefactor $F_{0}(b)$ depends only on the binary separation and is therefore fixed for a given lens mass distribution. In contrast, the final term, determined by the caustic width $d_{c}$, captures the frequency (or wavelength) dependence of the incoming wave. For EM waves, where the observations are based on intensity, the relevant quantity is the magnification $\mu=|F|^{2}=F_{0}^{2}\,d_{c}^{-1/2}$.
Therefore, the amplification $|F|_{\rm max}$ is governed by the characteristic scales of the caustic. Using Eq. (31), we find that it scales with the wavelength as $\sim\lambda^{-1/6}$, and with the lens mass as $\sim M_{Lz}^{1/6}$ as shown in Fig. 4. This leads to the following expression:
$|F|_{\rm max}\simeq 2.5\ |T_{11}|^{-1/2}\,\rho_{c}^{-1/3}\left(\frac{\lambda}{10^{6}{\rm m}}\right)^{-1/6}\left(\frac{M_{Lz}}{10^{3}M_{\odot}}\right)^{1/6}$ (35)
@@
As an example, consider a binary lens system with $b=0.7$ and a frequency of 500 Hz. In this case, we find that the maximum amplification is $|F|_{\rm max}\simeq 2$ for a lens mass of $M=5\times 10^{3}M_{\odot}$, and $|F|_{\rm max}\simeq 22.3$ for $M=10^{10}M_{\odot}$. For comparison, the corresponding magnification values for electromagnetic (EM) waves will be the square of these values.
@@ VI Conclusions
This study introduces a novel approach to understanding GW diffraction near caustics created by a binary lens system. The Uniform Approximation proves to be a robust method for modeling the GW field near the fold, addressing the limitations of both Geometrical Optics and Transitional Asymptotics methods. By ensuring continuity across the caustic, the UA effectively links wave behavior near the fold to the far-field GO description, providing a comprehensive solution for wave propagation in caustic regions. While the TA [Eq. (16)] and UA [Eqs. (27), (29)] are based solely on the generic structure of fold caustics and are therefore broadly applicable, the specific numerical predictions we present are tailored to the binary lens model and may differ quantitatively for other lenses producing fold caustics.
@@ Appendix A Images and caustics of a binary lens
where $\mathbf{x_{A}}$ and $\mathbf{x_{B}}$ are the positions of the two lensing masses on the lens plane, and $\gamma_{1}$ and $\gamma_{2}$ are the mass ratios defined in Sec. II. If the $x_{1}$ axis aligns with the line connecting the masses and the origin is at its midpoint, the mass coordinates satisfy $\mathbf{x_{A}}=-\mathbf{x_{B}}=\mathbf{x_{L}}=(b,0)$.
Figure 5: Source plane with caustics (left panel) and lens plane with critical lines (right panel) for a binary lens system with equal masses and $b=0.7$. A source is placed at two different positions, both at a distance of $|\tilde{y}_{2}|=0.05$ from the fold: one inside the caustic (red star) and one outside (black star). Their corresponding images in the lens plane are marked with dots of the same color. The location of the fold point is indicated by a black circle, while the point of coalescence of the two stationary points in the lens plane is denoted by a black triangle.
@@
Figure 8: Absolute value of the transmission factor computed with UA as a function of the frequency and the separation of the source with respect to the fold. The binary separation is fixed at $b=0.7$, and the lens mass is set to $M_{Lz}=5\times 10^{3}M\odot$. The fold location is indicated by a black line. The white continuous and dashed lines indicate the maxima computed with UA (40) and TA (41), respectively.
@@ Appendix B Interference Fringes
where $D=5\rho_{c}^{-1}\left[3/2+(2\rho_{c}^{\prime}/3)^{2}\right]^{-1}$. By substituting the phase difference by Eq. (39), we obtain the source positions corresponding to the extrema of the transmission factor. The dependence of these positions on the caustic curvature and the dimensionless frequency is nontrivial, resulting in a complex scaling for the fringe separation. Fig. 8 shows a colormap of the absolute value of the transmission factor as a function of the source’s distance from the fold and the frequency, where Eq. (40) is used to visualize the maxima and minima.